Find the Laplace transforms of the following function:

3. L{e^-5t + t^2}


1
Expert's answer
2022-05-22T17:45:31-0400

"L(af(t)+bg(t))=aL((f(t))+bL(g(t)))"

"L(e^{-5t}+t^2)=L(e^{-5t}+L(t^2))=\frac{1}{s+5}+\frac{2!}{s^3}"